# Constraint Learning for Parametric Point Cloud

### v1

@@ Abstract
Parametric point clouds are sampled from CAD shapes, have become increasingly prevalent in industrial manufacturing. However, most existing point cloud learning methods focus on the geometric features, such as local and global features or developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to fully comprehend CAD shapes. To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) is developed to extract and leverage constraints. Our CstNet includes two stages. The Stage 1 extracts constraints from B-Rep data or point cloud. The Stage 2 leverages coordinates and constraints to enhance the comprehend of CAD shapes. Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets. Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shapes datasets. To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shapes analysis.
@@ 1 Introduction
Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters. By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Figure 1. Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [32], PointNet++ [33], and PointCNN [24] demonstrating significant success across various tasks. These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours. However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions. This similarity in appearance poses challenge in distinguishing them based on geometric features. Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning. Thus, distinguishing them from constraints-based perspective becomes more feasible, as illustrated in Figure 1.
@@
Additionally, the limited availability of B-Rep datasets poses barrier to deep learning on CAD shapes. Most existing CAD shapes datasets consist of mesh files, such as MCB [16] and ESB [12]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data serves as the native representation of CAD shapes and is therefore more suitable for datasets construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development. Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relation. To address this challenge, we proposed a novel constraint representation as point-wise local attributes. Afterwards, we developed the CstNet for constraint extraction and feature learning, which consist of two stages. The stage 1 is built for constraint acquisition, if B-Rep data is avilable, the openCASCADE Technology (OCCT) is used to extracts constraints, otherwise the constraint prediction from point cloud (CST-PCD) module is adopted. The CST-PCD is designed relies on local features only, enabling it to generalize to unseen datasets after pre-training. The stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes. Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud. Given the current scarcity of labeled CAD shapes datasets, Parametric 20000 offers researchers an expanded resource for their work. Our contributions are threefold:
  • We introduce the deep learning friendly expression of constraints.
  • We designed the CstNet for constraint extraction and parametric feature learning.
  • We collected and curated the Parametric 20000 dataset, which consist of mesh, point cloud, and B-Rep data.
@@ 2 Related Work
For the published deep learning methods focus on CAD shape are very few, and even fewer aimed at parametric point cloud, this section will provide an overview of existing point cloud learning methods as well as CAD shapes oriented models. By integrating insights from above areas, a comprehensive review of the relevant works could be presented.
@@ 2.1 Deep learning on point clouds
Voxelization-based methods convert point cloud into 3D voxels, and then processed by 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20]. However, the size and orientation of the voxels can affect computational efficiency, memory usage. The voxel grid quantization process may lead to geometric detail lost, and the inherent sparsity of point clouds may result in many useless voxels.
Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 55, 56, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4]. Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction. However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
@@ 2.2 Deep learning on CAD shapes
B-Rep-based methods typically involve B-Rep data conversion and feature extraction. Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation is commonly used. These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 43], engineering sketch generation [11], and CAD operation sequence prediction [45, 57].
@@
Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22], or CAD operation sequences [51]. Most of these models are generative, and utilizing transformer modules frequently. A key challenge is incorporating the regularity and connectivity of primitives, which typically overlooked in graphic objects analysis.
@@ 3.1 Problem Statement
Constraints is crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance? To answer above question, we designed the following experiments. The objective is to compare the prisms shown in Figure 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, which predict constraints and then adopt it for classification, more details in supplementary.
@@
The experimental results are presented in Figure 3. For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results. The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid. For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced model’s comprehension on CAD shapes.
@@ 3.2 Constraint Representation
Following the constraint definition in CAD software is not applicable, However, once constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraint on CAD shape can be defined comprehensively.
The point-wise primitive type could be represented as the type of the primitive which the point aligned to, using one-hot encoding.
The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Figure 4, relations such as parallel and vertical could be derived by the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information.
@@
Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitation on the distance and size of other primitives. For example, in a cuboid, the adjacency and normal of the six faces ensures that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency combined with main axis directions, the positional relation can be determined comprehensively.
Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are the key indicator of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
In summary, each composition of constraints in parametric point cloud is illustrated in Figure 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
@@ 3.3 Method Overview
Our CstNet and module details are depicted in the top and bottom side of Figure 6 respectively. It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen. Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen mechanical datasets directly. In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
Figure 6: Overview of CstNet. The top side illustrates the overall architecture of CstNet, comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD predicts constraints. Stage 2 perform constraint feature learning, facilitate deeper understanding for CAD shapes. The bottom side presents the details of module design.
@@ 3.4 Constraint Acquisition
For adjacency computation, we identify valid edges firstly. In this study, an edge is considered invalid if the two adjacent faces are smoothly connected along the edge, this work is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which face it aligned with, OCCT is then used to analyze the face, providing both the face type and main axis direction.
@@
Step 2. From all neighbor points, identify which belong to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
Step 3. Fit shapes such as cylinders or planes using the valid points, and determine the primitive type with the smallest fitting error.
Step 4. Based on the the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
@@
The design of CST-PCD module is inspired by the aforementioned calculation process. For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Figure 7. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
@@
For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applying different weight to neighbor points when updating the center point’s embedding. We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Equation (1).
@@
The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset. As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve for meaningful tasks.
@@ 3.5 Constraint Feature Learning
The stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes. The designed architecture extract features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature. The structure and module details are illustrated in Figure 6.
@@
In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feed into MLP, with the output being the Parametric Feature.
@@
The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update constraint feature efficiently. This design is considered that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature. The Fea Attention is developed by vector attention as formulated in Equation (2).
@@ 4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Figure 8. By referring the model row, this module performs well in extracting three components of constraint.
Extract constraints from point cloud is achieved by CST-PCD module. The training set is sourced from 25 Trunks of ABC [18], with each trunk containing 10,000 B-Rep files. We utilized CST-BRep to extract constraints as Ground Truth for model training. After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
In subsequent classification and segmentation tasks, above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
After CST-PCD module is pre-trained, we evaluated it on a new Trunk of ABC. The visualization of constraint prediction is shown in the "Pred." column of Figure 8. After traversing this Trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
@@
Additionally, we evaluated the pre-trained CST-PCD module on the CAD shapes dataset MCB [16] to predict constraints. For the MCB consists of mesh files, it was not possible to extract ground truth constraints for comparison. Based on the results visualized in Figure 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.

### v2

@@ Abstract
Parametric point clouds are sampled from CAD shapes, and have become increasingly prevalent in industrial manufacturing. However, most existing point cloud learning methods focus on the geometric features, such as developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to comprehend CAD shapes fully. To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) was developed to extract and leverage constraints. Our CstNet includes two stages. Stage 1 extracts constraints from B-Rep data or point cloud. Stage 2 leverages coordinates and constraints to enhance the comprehension of CAD shapes. Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets. Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shape datasets. To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shape analysis.
@@ 1 Introduction
Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters. By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Fig. 1. Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [32], PointNet++ [33], and PointCNN [24] demonstrating significant success across various tasks. These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours. However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions. This similarity in appearance poses challenges in distinguishing them based on geometric features. Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning. Thus, distinguishing them from a constraint-based perspective becomes more feasible, as illustrated in Fig. 1.
@@
Additionally, the limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes. Most existing CAD shape datasets consist of mesh files, such as MCB [16] and ESB [12]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data [19, 56] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development. Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relations. To address this challenge, we proposed a novel constraint representation as point-wise local attributes. Afterward, we developed the CstNet for constraint extraction and feature learning, which consist of two stages. Stage 1 is built for constraint acquisition, if B-Rep data is available, the openCASCADE Technology (OCCT) is used to extract constraints, otherwise, the constraint prediction from point cloud (CST-PCD) module is adopted. The CST-PCD is designed to rely on local features only, enabling it to generalize to unseen datasets after pre-training. Stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes. Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud. Given the current scarcity of labeled CAD shape datasets, Parametric 20000 offers researchers an expanded resource for their work. Our contributions are threefold:
  • We introduced the deep learning-friendly expression of constraints.
  • We designed the CstNet for constraint extraction and constraint feature learning.
  • We collected and curated the Parametric 20000 dataset, which consists of mesh, point cloud, and B-Rep data.
@@ 2 Related Work
For the published deep learning methods focusing on CAD shapes are very few, and even fewer aimed at parametric point clouds, this section will provide an overview of existing point cloud learning methods as well as CAD shapes-oriented models. By integrating insights from the above areas, a comprehensive review of the relevant works could be presented.
@@ 2.1 Deep learning on point clouds
Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [30, 47, 35, 40, 17, 20]. However, the size and orientation of the voxels can affect computational efficiency and memory usage. The voxel grid quantization process may lead to geometric detail loss, and the inherent sparsity of point clouds may result in many useless voxels.
Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [32, 33, 24, 10, 52, 2, 38, 27, 58, 59, 9, 7, 50, 49, 48, 34, 28, 26, 41, 46, 4, 57]. Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction. However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
@@ 2.2 Deep learning on CAD shapes
B-Rep-based methods typically involve B-Rep data conversion and feature extraction. Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation commonly used. These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 43], engineering sketch generation [11], and CAD operation sequence prediction [45, 60].
@@
Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 44, 31, 36, 25], 3D models [21, 22, 55], or CAD operation sequences [51]. Most of these models are generative, and frequently utilize transformer modules. A key challenge is incorporating the regularity and connectivity of primitives, which are typically overlooked in graphic object analysis.
@@ 3.1 Problem Statement
Constraints are crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance? To answer above question, we designed the following validation experiments. The objective is to compare the prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, this model predicts constraints and then adopts it for classification. The constraint representation is defined as the relation between the point attached primitive and a reference plane, more details in supplementary.
@@
The experimental results are presented in Fig. 3. For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results. The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid. For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced the model’s comprehension of CAD shapes.
@@ 3.2 Constraint Representation
Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding.
The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information.
@@
Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively.
Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
In summary, each composition of constraints in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
@@ 3.3 Method Overview
Our CstNet and module details are depicted in the top and bottom sides of Fig. 6 respectively. It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen. Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen CAD shape datasets directly. In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
Figure 6: Overview of CstNet. The top side illustrates the overall architecture of CstNet, comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design.
@@ 3.4 Constraint Acquisition
For adjacency computation, we identify valid edges first. In this study, an edge is considered invalid if the two adjacent primitives are smoothly connected along the edge, this work is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
@@
Step 2. From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
Step 3. Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
Step 4. Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
@@
The design of CST-PCD module is inspired by the aforementioned calculation process. For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 7. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
@@
For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding. We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Eq. 1.
@@
The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset. As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
@@ 3.5 Constraint Feature Learning
Stage 2 of CstNet is designed to leverage constraints to deepen the model’s understanding of CAD shapes. The designed architecture extracts features from Primitive Type, Main Axis Direction, and Adjacency separately, attention between features is then applied to update constraint feature. The structure and module details are illustrated in Fig. 6.
@@
In the C-MLP layer, point cloud is concatenated with Primitive Type, Main Axis Direction, Adjacency, and then feeds into MLP, with the output being the Parametric Feature.
@@
The Fea Attention layer also takes the Primitive Feature, Axis Feature, Adjacency Feature, and Parametric Feature as input, attention between features is employed to update the constraint feature efficiently. This design considers that different points may focus on different features, for example, points near edge may focus more on the Adjacency Feature, while others may pay more attention to the Primitive or Axis Feature. The Fea Attention is developed by vector attention as formulated in Eq. 2.
@@ 4.2 Constraint Acquisition
The extraction of constraints from B-Rep data is accomplished by CST-BRep module, with results visualized in the "GT" column of Fig. 8. By referring to the model row, this module performs well in extracting three components of constraint.
Extracting constraints from point cloud is achieved by CST-PCD module. The training set is sourced from 25 trunks of ABC [18], with each trunk containing 10,000 B-Rep files. We utilized CST-BRep module to extract constraints as Ground Truth for model training. After the training finished, the prediction accuracy of Primitive Type reached 94.23 %, for Adjacency, that is 94.62 %, and the MSE loss for Main Axis Direction is 0.0559.
In subsequent classification and segmentation tasks, the above pre-trained CST-PCD module is utilized to predict constraints with its weights frozen and no further updates. To ensure that only point cloud is used as input across all methods in comparison, even when B-Rep data is available, the CST-BRep module will not be activated to obtain constraints.
After CST-PCD module was pre-trained, we evaluated it on a new trunk of ABC. The visualization of constraint prediction is shown in the "Pred." column of Fig. 8. After traversing this trunk, the prediction accuracy of Primitive Type reached 91.19 %, for Adjacency, that is 90.24 %, the MSE loss of Main Axis Direction is 0.0998, above demonstrating the strong capability for constraint prediction.
@@
Additionally, we evaluated the pre-trained CST-PCD module on the CAD shape dataset MCB [16] to predict constraints. For the MCB consists of mesh files, it was not possible to extract Ground Truth constraints for comparison. Based on the results visualized in Fig. 9, CST-PCD demonstrated a high level of constraint prediction accuracy even on an unseen dataset.